A car accelerate for 50 seconds at a starting 100 kilometers per hour. what is the velocity of the car at end time of 100 seconds if it has an acceleration of 0.6 meter per seconds squared?
1
Expert's answer
2018-02-26T09:06:07-0500
100 km/h = 100/3.6 = 27.8 m/s
v = v0 + at = 27.8 + 0.6x100 = 87.8 m/s = 316 km/h
Answer: 316 km/h or 87.8 m/s.